### v1

@@ 6 Performance Overview
UnlabeledHybrid VideoMAEv2 [Wang et al. [41]] 86.9
ViT-G UnlabeledHybrid VideoMAEv2 [Wang et al. [41]] 87.2
IN21K EVA [Fang et al. [63]] 89.7
@@
MViTv2-S K400 MaskFeat [Wei et al. [171]] 82.2
MViTv2-L K400 MaskFeat [Wei et al. [171]] 86.7
@@ 7 Closing Remarks and Future Directions
While many papers focused on images as the only input modality, some promising results were shown on multimodal inputs. Each modality holds a different type of information, and these can be combined to learn richer feature representations. Distilling the knowledge of each modality source into a single network could represent a stepping stone in artificial intelligence.

### v2

@@ 6 Performance Overview
UnlabeledHybrid VideoMAEv2 [57] 86.9
ViT-G UnlabeledHybrid VideoMAEv2 [57] 87.2
IN21K EVA [75] 89.7
@@
MViTv2-S K400 MaskFeat [73] 82.2
MViTv2-L K400 MaskFeat [73] 86.7
Other significant observations, based on the results presented in Table 4, refer to the masking strategy and the types of features that are reconstructed. First, in videos, the temporal correlation between frames [35] can introduce a bias in the model, allowing it to reconstruct masked patches by leveraging information from adjacent frames. To alleviate this issue, OmniMAE [38] randomly masks patches from the videos similar to MAE [9], but uses a higher masking ratio, $95\%$ instead of $75\%$. However, this masking strategy performs poorly compared with those used in VideoMAE [35] and VideoMAEv2 [57], which are designed to mask temporal tubes instead of spatial patches. The latter approach is more effective in preventing information leakage from adjacent frames. Second, the top performing methods, such as MVD [183] and EVA [75], benefit from the use of high-level features as target features for reconstruction. These features help address the tendency of pixel-level MIM models to focus on spatial and structural information, leading to improved performance on downstream tasks.
@@ 7 Closing Remarks and Future Directions
While many papers focused on images as the only input modality, some promising results were shown on multimodal inputs. Each modality holds a different type of information, and these can be combined to learn richer feature representations. Distilling the knowledge of each modality into a single network could represent a stepping stone in artificial intelligence.
Declarations
Funding. This research is supported by the project “Romanian Hub for Artificial Intelligence - HRIA”, Smart Growth, Digitization and Financial Instruments Program, 2021-2027, MySMIS no. 334906.
Conflict of interest. The authors have no conflicts of interest to declare that are relevant to the content of this article.
Availability of data and materials. This paper is a survey paper, hence it does not report new experiments on any dataset.